# Machine Translation in the Wild: User Reaction to Xiaohongshu’s Built-In Translation Feature

### 16 Mar 2026 v1

@@ 5.3 Translation in a Prominent Socio-Technical System
User comments also reveal that perceptions of the translation feature extend beyond the quality of the translation output itself. Many users commented on the speed with which the function was introduced and expanded across different parts of the platform. Some expressed admiration for the rapid development cycle, while others voiced concern about the workload placed on developers and programmers. Such remarks indicate that users are aware of the human actors involved in producing and maintaining the system, even when interacting with automated functions.
@@
These observations highlight the inherently social nature of translation technologies in digital environments. MT on social platforms emerges through the interaction of multiple stakeholders, including users, developers, designers, translators, and language experts. Each group contributes to how translation functions are designed, implemented, interpreted, and evaluated. Consequently, assessments of translation quality in contexts as such should move beyond laboratory-style benchmarks and controlled metrics. While these measures remain important for comparing system performance and tracking technical improvements, they capture only part of how translation systems function in practice.
Real-world deployments of translation features generate large-scale user interactions that provide valuable insight into usability, perceived reliability, and communicative effectiveness. User experimentation, feedback, and informal testing collectively form a form of situated evaluation that reflects how translation technologies operate within everyday communication. Incorporating these perspectives into research and development may therefore contribute to more user-centred approaches to MT design and evaluation, ensuring that technological advancements are aligned with the communicative needs and expectations of the communities that use them.
@@ 6 Conclusion
This study examined early user responses to the introduction of Xiaohongshu’s built-in translation feature by combining sentiment analysis with thematic analysis of user comments and test inputs. The findings suggest that user perceptions of translation functions are shaped by a range of socio-technical factors, including testing practices, feature accessibility, and users’ awareness of the processes involved in system development. In this sense, translation features on social media platforms are experienced not merely as technical outputs but as components of a broader communicative ecosystem involving users, developers, and language professionals.
The analysis also indicates that the types of text segments users selected to test the system were relatively limited. As a result, users’ evaluations of translation performance may reflect a narrow range of use cases, potentially shaping perceptions of the function’s usefulness in everyday cross-lingual communication. At the same time, the findings point to the need for greater public awareness of what constitutes an appropriate translation, particularly in contexts where automated translation tools are increasingly integrated into routine online interactions.
More broadly, incorporating user-generated feedback and interaction patterns into translation research can contribute to a more comprehensive understanding of how LLM-powered translation tools function in real-world multilingual communication. Such perspectives can shed light on how these systems are perceived, trusted, adopted, and shaped by users in practice.
Admittedly, this study provides only a snapshot of how users responded to an LLM-enabled MT feature when it was first introduced on Xiaohongshu. Further research is needed to develop a more comprehensive understanding of how the function is used in real-time communication beyond the context of the 11 posts analysed here. Nevertheless, it is clear that addressing the challenges of deploying MT systems on social media platforms will require closer interdisciplinary collaboration, bringing together computer scientists, translation scholars, and other stakeholders. Improving MT performance in practice ultimately depends not only on advances in algorithms, but also on sustained dialogue between those who build these systems and those who study and use them.

### 03 Apr 2026 v2

@@ 5.3 Translation in a Socio-Technical System
User comments also reveal that perceptions of the translation feature extend beyond the quality of the translation output itself. Some expressed admiration for the rapid development cycle, while others voiced concern about the workload placed on developers and programmers. Such remarks indicate that users are aware of the humans involved in producing and maintaining the system, even when interacting with automated functions.
@@
These observations highlight the inherently social nature of translation technologies in digital environments. MT on social platforms emerges through the interaction of multiple stakeholders, including users and marketing specialists at the forefront and MT specialists and linguists behind the architecture, in a situated socio-economic context. Each group of stakeholders contributes to how translation functions are designed, implemented, interpreted, and evaluated. Consequently, in a real-life context as such, the assessments of translation quality should move beyond laboratory-style benchmarks. It is worth reiterating that while these measures remain important for comparing system performance and tracking technical improvements, they capture only part of how translation systems function in practice. This is particularly important because even though MT outputs can contain errors and impose risks across communicative contexts, its convenience can in fact ‘shape perceptions and expectations of what translations should be and the different roles they can play’ [20].
@@ 6 Conclusion
This study examined early user responses to the introduction of Xiaohongshu’s built-in MT feature. The findings suggest that user perceptions of translation functions are shaped by a range of socio-technical factors, including promotional materials of the tools, testing practices, and feature accessibility. In this sense, translation features on social media platforms are experienced not merely as technical outputs but as components of a broader communicative ecosystem. The analysis also indicates that the types of text segments users selected to test the system were relatively limited. As a result, users’ evaluations of MT performance may reflect a narrow range of use cases, potentially shaping perceptions of the function’s usefulness in everyday cross-lingual communication, calling for better MT literacy among the general public, particularly in contexts where MT tools are increasingly integrated into daily interactions.
More broadly, real-world deployments of translation features generate large-scale user interactions that provide valuable insight into usability, perceived reliability, and communicative effectiveness. User experimentation, feedback, and informal testing collectively form a form of situated evaluation that reflects how translation technologies operate within everyday communication. Incorporating these perspectives into research and development may therefore contribute to more user-centred approaches to MT design and evaluation, ensuring that technological advancements are aligned with the communicative needs and expectations of the communities that use them.
Admittedly, this study provides only a snapshot of how users responded to an LLM-enabled MT feature when it was first introduced on Xiaohongshu. Further research is needed to develop a more comprehensive understanding of how the function is used in real-time communication beyond the context of user engagement with the 11 official posts analysed here. Nevertheless, it is clear that addressing the challenges of deploying MT systems on social media platforms will require closer interdisciplinary collaboration, bringing together computer scientists, translation scholars, users, and other stakeholders. Improving MT performance in practice ultimately depends not only on advances in algorithms, but also on sustained dialogue between those who build these systems and those who study and use them. As Kenny [13] observes, ‘it is not the technology alone that shapes the future; rather it is the way in which it is accommodated by the socio-cultural, legal and economic context, itself shifting in line with technological change that will have the greatest bearing’.